### v1

# A TYPE $\textrm{III}_{1}$ FACTOR WITH THE SMALLEST OUTER AUTOMORPHISM GROUP

Abstract.
The canonical modular homomorphism provides an embedding of $\mathbb{R}$ into the outer automorphism group $\operatorname{Out}(M)$ of any type $\rm{III}_{1}$ factor $M$. We provide an explicit construction of a full factor of type $\rm{III}_{1}$ with separable predual such that the outer automorphism group is minimal, i.e. this embedding is an isomorphism. We obtain such a $\rm{III}_{1}$ factor by using an amalgamated free product construction.
1. Introduction
In this article we will only deal with von Neumann algebras with separable preduals. An important invariant for the study of $\rm{II}_{1}$ factors is the outer automorphism group of the factor. Recall that an automorphism of such a factor $M$ is called inner if it is of the form $\operatorname{Ad}(u)$ for a unitary $u\in M$. The outer automorphism group $\operatorname{Out}(M)$ is defined as the quotient of the group of all automorphisms of $M$ by the group of inner automorphisms. In [Bla58], Blattner showed that $\operatorname{Out}(R)$ for the hyperfinite $\rm{II}_{1}$ factor $R$ is huge: it contains an isomorphic copy of every locally compact second countable group. In general, computation of the outer automorphism group for a $\rm{II}_{1}$ factor is a difficult problem. The first significant step in this direction was taken in [Con80] where Connes introduced the notion of property (T) von Neumann algebras and showed that any property (T) factor has countable outer automorphism group. In particular he showed that group von Neumann algebras $L(G)$ of property (T) groups $G$ with infinite conjugacy classes (ICC) have countable outer automorphism groups. An interesting and natural question then was if there exists a $\rm{II}_{1}$ factor $M$ with trivial outer automorphism group. This was a difficult question and remained unanswered for many years until Popa’s deformation and rigidity theory made it possible to attack such problems.
In their seminal article [IPP08], Ioana, Peterson and Popa showed that any second countable compact abelian group appears as the outer automorphism group of a $\rm{II}_{1}$ factor. In [FV08], Falguières and Vaes improved this result and showed that in fact every compact group (not necessarily abelian) arises as outer automorphism group of a $\rm{II}_{1}$ factor. However all these results used a Baire category argument which showed the existence of such $\rm{II}_{1}$ factors, but did not provide explicit constructions. In [PV10], Popa and Vaes gave explicit constructions of certain $\rm{II}_{1}$ factors with prescribed outer automorphism groups, including the first explicit example of a $\rm{II}_{1}$ factor with trivial outer automorphism group. It is still unknown if every locally compact unimodular group arises as the outer automorphim group of a $\rm{II}_{1}$ factor. However by now we have a large class of examples of such groups: in [Vae08], Vaes showed that every discrete countable group arises as an outer automorphism group, we earlier saw that every compact group arises this way too. In [Dep12], Deprez showed that the groups $\operatorname{SL}^{\pm}(n,\mathbb{R})$ also arise as outer automorphism groups of $\rm{II}_{1}$ factors.
This article deals with the type $\rm{III}$ version of this problem. If $M$ is a type $\rm{III}$ factor with a faithful normal state $\phi$, then the modular automorphism group $(\sigma^{\phi}_{t})_{t\in\mathbb{R}}$ form a natural class of automorphisms of $M$. A deep result of Connes showed that the modular automorphism group does not depend on the choice of the faithful normal state up to inner automorphisms. If furthermore $M$ is of type $\rm{III}_{1}$, then the modular automorphisms are outer, and hence we have that $\mathbb{R}\subseteq\operatorname{Out}(M)$. We are interested in the question: can we explicitly construct a factor $M$ of type $\rm{III}_{1}$ such that $\operatorname{Out}(M)=\mathbb{R}$. We tackle this problem with ergodic theoretic methods: by using group measure space von Neumann algebras, where the outer automorphism group has a more tractable description in terms of the outer automorphism group of the associated equivalence relation, as below.
Suppose that we have a free ergodic nonsingular action $\Gamma\curvearrowright(Z_{1},\eta_{1})$ for a discrete countable group $\Gamma$ and let the associated countable ergodic equivalence relation on $(Z_{1},\eta_{1})$ be $\mathcal{R}_{1}$ (the subscript 1 is used intentionally as we deal with actions of type $\rm{III}_{1}$). Recall that an automorphism of $\mathcal{R}_{1}$ is a nonsingular Borel isomorphism $\theta:Z_{1}\rightarrow Z_{1}$ such that for a.e. $z,z^{\prime}\in Z_{1}$, $(z,z^{\prime})\in\mathcal{R}_{1}$ if and only if $(\theta(z),\theta(z^{\prime}))\in\mathcal{R}_{1}$. Such an automorphism is called inner if $(z,\theta(z))\in\mathcal{R}_{1}$ for a.e. $z\in Z_{1}$. The outer automorphism group $\operatorname{Out}(\mathcal{R}_{1})$ is defined similarly as the quotient of the automorphism group by the inner ones. By the theory of Feldman and Moore ([FM77], [FM77a]), denoting by $M_{1}$ the von Neumann algebra $L^{\infty}(Z_{1},\eta_{1})\rtimes\Gamma$ and by $A$ the Cartan subalgebra $L^{\infty}(Z_{1},\eta_{1})$, we have that:
(1) $\operatorname{Out}(A\subset M_{1})=\operatorname{Out}(\mathcal{R}_{1})\ltimes H^{1}(\mathcal{R}_{1},S^{1})$
where $\operatorname{Out}(A\subset M_{1})$ denotes the group of outer automorphisms preserving the Cartan subalgebra $A$ and $H^{1}(\mathcal{R}_{1},S^{1})$ denotes the quotient of the 1-cocycle group of $\mathcal{R}_{1}$ with values in the circle, by the group of coboundaries. In particular if $A\subset M_{1}$ is the unique Cartan subalgebra up to unitary conjugacy, then equation 1 indeed becomes:
(2) $\operatorname{Out}(M_{1})=\operatorname{Out}(\mathcal{R}_{1})\ltimes H^{1}(\mathcal{R}_{1},S^{1})$
The modular automorphism group sits inside the 1-cohomolgy group $H^{1}(\mathcal{R}_{1},S^{1})$. Therefore the idea is to construct a ‘rigid’ free ergodic action $\Gamma\curvearrowright(Z_{1},\eta_{1})$ of type $\rm{III}_{1}$ such that every orbit equivalence is inner, the 1-cohomology group with values in the circle is equal to $\mathbb{R}$ and the associated von Neumann algebra has a unique Cartan subalgebra up to unitary conjugacy. In [Ioa15], Ioana showed that for a large class of amalgamated free product groups, any free ergodic pmp action has the property that the crossed product has a unique Cartan subalgebra. Notice that if $\Gamma=\Gamma_{1}\ast_{\Sigma}\Gamma_{2}$ is an amalgamted free product group and $\Gamma\curvearrowright(Z,\eta)$ is a nonsingular free ergodic action which is induced from an action $\Gamma_{i}\curvearrowright(X,\mu)$, then $L^{\infty}(Z)\rtimes\Gamma$ is isomporphic to $B(\mathcal{H})\overline{\otimes}(L^{\infty}(X)\rtimes\Gamma_{i})$, and in such a situation we cannot have uniqueness of Cartan subalgebras. However this can be taken care of by imposing some recurrence of $\Gamma_{i}\curvearrowright(Z,\eta)$ relative to $\Sigma$. Using the methods of [HV13], Vaes in [Vae14, Theorem 8.1] proved that under some suitable recurrence assumptions, any non singular free ergodic action has the property that the crossed product has a unique Cartan subalgebra, generalizing the unique Cartan decomposition results in [BHR14]. We use a slightly modified version of this result that we state in Theorem 5.5, which was already observed in [Vae14, Remark 8.3].
In this article we construct such an action of an amalgamated free product group that is sufficiently rigid. Orbit equivalence (OE)-superrigidity results for an action usually follow from cocycle superrigidity with countable target groups. Over the years a lot of these results have been proven for example for Bernoulli actions of property (T) groups in [Pop07], of product groups in [Pop08] and for profinite actions of property (T) groups in [Ioa11]. In [PV11] the authors use similar methods to prove cocycle and OE-superrigidity results for actions of lattices in $\operatorname{SL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$. Since these actions are not probability measure preserving any more, the notion of property (T) has to be replaced by Zimmer’s notion of property (T) for non singular actions, and this plays a very important role in [PV11]. Using such an action of a lattice and the unique Cartan results, Vaes in [Vae14, Proposition D] constructed the first example of a type III action which is W*-superrigid. In [VV23, Remark 6.10], the authors used this construction to give examples of type $\rm{III}_{0}$ actions that are W*-superrigid with any prescribed flow. Our construction in this article is similar to the one in [Vae14, Proposition D] but the main difficulty is that we need to make suitable modifications to take care of $\operatorname{Out}(\mathcal{R}_{1})$ and the 1-cohomology group $H^{1}(\mathcal{R}_{1},S^{1})$.
Now we explain our construction. We take two countable discrete groups $\Gamma_{1}$ and $\Lambda$ and let $\Sigma<\Gamma_{1}$ be a proper subgroup. We consider the amalgamated free product $\Gamma=\Gamma_{1}*_{\Sigma}(\Sigma\times\Lambda)$ with the natural quotient morphism $\pi:\Gamma\rightarrow\Gamma_{1}$. We will consider two actions: a free ergodic probability measure preserving action of $\Gamma\curvearrowright(X,\mu)$ and a free ergodic nonsingular action $\Gamma_{1}\curvearrowright(Y_{1},\nu_{1})$ of type $\textrm{III}_{1}$ such that the infinite measure preserving Maharam extension $\Gamma_{1}\curvearrowright(Y,\nu)$ is ergodic. Putting $(Z_{1},\eta)=(X\times Y_{1},\mu\times\nu_{1})$, we show that the diagonal action $\Gamma\curvearrowright Z_{1}$ given by $g\cdot(x,y)=(gx,\pi(g)\cdot y)$ satisfies the required properties. We take $\Gamma_{1}\curvearrowright Y_{1}$ to be the rigid lattice action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ such that the Maharam extension $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}$ is ergodic. We then pick a subgroup $H<\Gamma$ in a way such that the conjugates of $\Sigma$ does not intersect $H$ except at ±$\pm{I}$ and take $\Gamma\curvearrowright X$ to be the generalized Bernoulli action $\Gamma\curvearrowright\{0,1\}^{\Gamma/H}$ with an unequal measure on the atomic base space $\{0,1\}$. In Section 4 we show that the Maharam extension of this action is cocycle superrigid with any target group in Popa’s class of $\mathcal{U}_{\operatorname{fin}}$ groups (see Definition 4.1). We essentially show that any 1 cocycle restricted to $\operatorname{SL}(6,\mathbb{Z})$ is cohomologous to a group homomorphism using [PV11, Theorem 5.3]. The reason we choose $H$ to not intersect conjugates of $\Sigma$ is so that the action $\Sigma\curvearrowright X$ remains mixing, which we need to show cocycle superrigidity. Finally we use the fact that the group $\Gamma$ is perfect to deduce that the 1 cohomology group is isomorphic to $\mathbb{R}$.
In Section 5 we finally compute the outer automorphism group. We apply cocycle superrigidity to the Zimmer cocycle of an orbit equivalence $\Delta$ of the Maharam extension. We use a general cocycle superrigidity to OE-superrigidity result proven in [DV23, Lemma 2.4] to see that $\Delta(g\cdot z)=\delta(g)\cdot\Delta(z)$ for an automorphism $\delta\in\operatorname{Aut}(\Gamma)$. We then use some properties of automorphisms of amalgamated free product groups as in Lemma 2.12 and Lemma 2.13 to deduce that $\delta$ is trivial and $\Delta$ is indeed a conjugacy. In Proposition 5.4 we note that $\Delta$ is trivial on the generalized Bernoulli action and on $\mathbb{R}^{6}$, we have that $\Delta(y)=g\cdot y$ for some some $g$ normalizing $\operatorname{SL}(6,\mathbb{Z})$ in $\operatorname{GL}(6,\mathbb{R})$ as in [PV11, Theorem 6.2]. Thus we deduce that $\Delta$ is of the form $(\operatorname{id}\times\rho_{\lambda})$ where $\rho_{\lambda}(y)=\lambda y$ for a non zero real number $\lambda$ in $\mathbb{R}^{6}$. Finally we note in Theorem 5.8 that the induced automorphism of $\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ is inner. Together with the fact that the 1 cohomology group is equal to $\mathbb{R}$ and that the Cartan subalgebra $L^{\infty}(Z_{1},\eta_{1})$ is unique up to unitary conjugacy, we have the required $\rm{III}_{1}$ factor with the smallest outer automorphism group.
Acknowledgements
The author is supported by FWO research project G090420N of the Research Foundation Flanders. The author would like to thank his PhD supervisor Stefaan Vaes for his immense support throughout the process of writing this article. The author would also like to thank Bram Verjans for a lot of helpful discussions on the topic.
2. Preliminaries
2.1. Facts about amalgamated free products
We begin this section with the following definition:
Definition 2.1.
Let $\Gamma_{1}=\operatorname{SL}(6,\mathbb{Z})$ and let $\Sigma$ be the subgroup of $\operatorname{SL}(6,\mathbb{Z})$ given by
(3) $\Sigma=\left(\begin{array}[]{@{}c|c@{}}\pm 1&\mathbb{Z}^{1\times 5}\\ \hline\cr 0&\operatorname{GL}(5,\mathbb{Z})\end{array}\right)\cap\operatorname{SL}(6,\mathbb{Z})$
We put $\Lambda=\operatorname{SL}(3,\mathbb{Z})$ and consider the amalgamated free product group $\Gamma=\Gamma_{1}*_{\Sigma}(\Sigma\times\Lambda)$ with the natural quotient map $\pi:\Gamma\rightarrow\Gamma_{1}$.
We shall now state several results about these groups that we need to prove our main theorem. We begin with the following well-known result about automorphisms of linear groups. For a proof of this fact, we point the reader to [HR52].
Proposition 2.2.
Let $G=\operatorname{SL}(n,\mathbb{Z})$ for $n\geq 3$. Then the only outer automorphisms of $G$ are as follows:
  1. (1)
    If $n$ is odd, $|\operatorname{Out}(G)|=2$, the only nontrivial one being $\phi(A)=(A^{T})^{-1}$.
  2. (2)
    If $n$ is even, $|\operatorname{Out}(G)|=4$ and is generated by the $\phi$ as above and $A\mapsto BAB^{-1}$ for an element $B\in\operatorname{GL}(n,\mathbb{Z})$ with determinant -1.
Definition 2.3.
Let $\Sigma<\operatorname{SL}(6,\mathbb{Z})$ as in Definition 2.1. We define the subgroup $\widetilde{\Sigma}$ of $\operatorname{GL}(6,\mathbb{Z})$ as $\widetilde{\Sigma}=\{g\in\operatorname{GL}(6,\mathbb{Z})\;|\;g(e_{1})=\pm e_{1}\}$. Note that $\widetilde{\Sigma}$ can be described as follows:
(4) $\widetilde{\Sigma}=\left(\begin{array}[]{@{}c|c@{}}\pm 1&\mathbb{Z}^{1\times 5}\\ \hline\cr 0&\operatorname{GL}(5,\mathbb{Z})\end{array}\right)$
Proposition 2.4.
Any automorphism $\phi$ of $\operatorname{SL}(6,\mathbb{Z})$ such that $\phi(\Sigma)=\Sigma$ is of the form $\phi=\operatorname{Ad}(B)$ with $B\in\widetilde{\Sigma}$.
Proof.
First we note the following: The globally invariant proper subspaces of $\mathbb{R}^{6}$ under the subgroup $\Sigma^{T}$ has to be of dimensions 5. On the other hand the globally invariant proper subspaces of $\mathbb{R}^{6}$ under $\Sigma$ must have dimension 1. Since conjugation does not change the dimension of a globally invariant subspace, we have that $\Sigma^{T}$ cannot be conjugated onto $\Sigma$. By Proposition 2.2, we have that such an automorphism has to be of the form $A\mapsto BAB^{-1}$ with $B\in\operatorname{GL}(6,\mathbb{Z})$ and $\det(B)=\pm 1$. If $W\subset\mathbb{R}^{6}$ is a one dimensional globally invariant subspace for $\Sigma$, then $B\cdot W$ is a globally invariant subspace for $B\Sigma B^{-1}$. Thus if $B\Sigma B^{-1}=\Sigma$ for some $B\in\operatorname{GL}(6,\mathbb{Z})$ then $B\cdot W=W$ and this forces $B$ to be in $\widetilde{\Sigma}$. ∎
The rest of this section will be dedicated to constructing a suitable subgroup $H$ of $\Gamma$ as in Definition 3.1.
Definition 2.5.
Let $p_{0}=5$, $p_{1}=7$ and $p_{2},...,p_{k}\geq 11$ be distinct primes. We define now the following elements of $\operatorname{SL}(3,\mathbb{Z})$ for $i=2,...,k$
(5) $A_{0}=\begin{pmatrix}1&0&0\\ 0&p_{0}-2&1\\ 0&-1&0\end{pmatrix},\text{ }A_{1}=\begin{pmatrix}p_{1}-2&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix},\text{ }A_{i}=\begin{pmatrix}p_{i}-2&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix}$
Lemma 2.6.
[…]
Proof.
Solving the characteristic polynomials, we note that the matrices $A_{i}$ have three distinct eigenvalues: 1, $\frac{p_{i}-2\pm\sqrt{p_{i}(p_{i}-4)}}{2}$. Since the $p_{i}$’s are distinct, two of these eigenvalues are nontrivial algebraic integers in distinct degree 2 field extensions of $\mathbb{Q}$. Denoting by $\operatorname{Spec}(A)$ the set of eigenvalues of a matrix $A$, we note the following:
  • $\operatorname{Spec}(A_{i})=\operatorname{Spec}(A_{i}^{-1})$ for all $i$
  • If $i\neq j$, for any nonzero integer $k_{i}$ and any integer $k_{j}$, we have that $\operatorname{Spec}(A_{i}^{k_{i}})\neq\operatorname{Spec}(A_{j}^{k_{j}})$
  • For a fixed $i$, if $k\neq l$, we have that $\operatorname{Spec}(A_{i}^{k})\neq\operatorname{Spec}(A_{i}^{l})$.
By Proposition 2.2, we have that $\sigma$ is either of the form $\operatorname{Ad}(Y)$ for $Y\in\operatorname{SL}(3,\mathbb{Z})$ or of the form $\operatorname{Ad}(Y)\circ\phi$ where $\phi(A)=(A^{T})^{-1}$. In particular we see the $\operatorname{Spec}(A)=\operatorname{Spec}(\sigma(A))$ for all $A\in\operatorname{SL}(3,\mathbb{Z})$. By the previous paragraph, we have hence that $\sigma(A_{i})=A_{i}^{\pm 1}$ for all $i$. Now suppose first that $\sigma=\operatorname{Ad}(Y)\circ\phi$ and consider the matrix
(6) $B=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix}$
and notice that for all $i\geq 2$, we have $\phi(A_{i})=BA_{i}B^{-1}$. Then letting $C=YB$, we have that $\sigma(A_{i})=CA_{i}C^{-1}$. Thus for all $i\geq 1$, $C$ either commutes with $A_{i}$ or $CA_{i}C^{-1}=A_{i}^{-1}$. Thus $C$ preserves globally the 1-dimensional eigenspace corresponding to the eigenvalue 1 and also preserves the 2-dimensional eigenspace corresponding to the two other eigenvalues of $A_{i}$. Thus $C$ has a simple block decomposition into a $2\times 2$ block and a $1\times 1$ block and by multiplying with $B^{-1}$, we know that $Y$ has the following form:
(7) $Y=\begin{pmatrix}*&*&0\\ *&*&0\\ 0&0&\pm 1\end{pmatrix}$
Now since $C$ also either commutes with $A_{0}$ or $CA_{0}C^{-1}=A_{0}^{-1}$, by the same argument we have that $Y$ is also necessarily of the following form:
(8) $Y=\begin{pmatrix}\pm 1&0&0\\ 0&*&*\\ 0&*&*\end{pmatrix}$
From the two equations of $Y$, we can conclude that $Y$ is a diagonal matrix with all entries equal to ±$\pm 1$ in the diagonal. But then we have that $Y(A_{0}^{T})^{-1}Y^{-1}\neq A_{0}^{\pm 1}$, hence giving us a contradiction. Thus $\sigma$ cannot be of the form $\operatorname{Ad}(Y)\circ\phi$. Suppose now that $\sigma=\operatorname{Ad}(Y)$. By a similar argument as earlier, we get once again that $Y$ is a diagonal matrix with ±$\pm 1$’s in the diagonal. Let us denote the diagonal entries of $Y$ by $\epsilon_{i}$ for $i=1,2,3$. For $i\geq 2$ we can then check that $YA_{i}Y^{-1}$ has to be equal to $A_{i}$ and this forces $\epsilon_{1}=\epsilon_{2}$. Similarly we can calculate $YA_{1}Y^{-1}$ and check that it is necessarily equal to $A_{1}$ and this forces $\epsilon_{1}=\epsilon_{3}$. Since $Y$ has determinant 1, this implies that $Y=\operatorname{id}$. ∎
Lemma 2.7.
For $n\geq 3$, consider the subgroup $\Sigma$ of $\operatorname{SL}(n,\mathbb{Z})$ given by $\Sigma=\{g\in\operatorname{SL}(n,\mathbb{Z})\;|\;g(e_{1})=\pm e_{1}\}$. As before consider the subgroup $\widetilde{\Sigma}$ of $\operatorname{GL}(n,\mathbb{Z})$ as $\widetilde{\Sigma}=\{g\in\operatorname{GL}(n,\mathbb{Z})\;|\;g(e_{1})=\pm e_{1}\}$. There exist elements $g_{1},...,g_{r}\in\operatorname{SL}(n,\mathbb{Z})\backslash\Sigma$ such that
(9) $\bigcap_{i=1}^{r}g_{i}^{-1}\widetilde{\Sigma}g_{i}=\{\pm I\}$
Moreover the $g_{i}$’s can be chosen such thath $g_{i}g_{j}^{-1}\notin\Sigma$ if $i\neq j$.
Proof.
For $2\leq i\leq n$, we define the elements $\sigma_{i}$ and $\alpha_{i}$ of $\operatorname{SL}(n,\mathbb{Z})\backslash\Sigma$ as the following linear operators:
(10) $\sigma_{i}:\begin{cases}e_{1}\mapsto-e_{i}\\ e_{i}\mapsto e_{1}\\ e_{j}\mapsto e_{j}\text{ for }j\neq\{1,i\}\end{cases}$
(11) $\alpha_{i}:\begin{cases}e_{1}\mapsto e_{1}+2e_{i}\\ e_{i}\mapsto e_{1}+3e_{i}\\ e_{j}\mapsto e_{j}\text{ for }j\neq\{1,i\}\end{cases}$
Now restricted to the linear subspace $\mathbb{C}e_{1}+\mathbb{C}e_{i}$, the matrices for the operators $\sigma_{i}$ and $\alpha_{i}$ are given by $\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$ and $\left(\begin{array}[]{cc}1&1\\ 2&3\end{array}\right)$ respectively. We claim now that
(12) $(\bigcap_{i=2}^{n}\sigma_{i}^{-1}\widetilde{\Sigma}\sigma_{i})\cap(\bigcap_{j=2}^{n}\alpha_{j}^{-1}\widetilde{\Sigma}\alpha_{j})=\{\pm I\}$
Clearly we have that $\sigma_{i}^{-1}\widetilde{\Sigma}\sigma_{i}=\{g\in\operatorname{GL}(n,\mathbb{Z})\;|\;g(e_{i})=\pm e_{i}\}$. Consider now an element $A\in(\bigcap_{i=2}^{n}\sigma_{i}^{-1}\widetilde{\Sigma}\sigma_{i})\cap(\bigcap_{j=2}^{n}\alpha_{j}^{-1}\widetilde{\Sigma}\alpha_{j})$. In particular, $A\in\sigma_{i}^{-1}\widetilde{\Sigma}\sigma_{i}$ which then forces $A(e_{i})=\pm e_{i}$ for all $2\leq i\leq n$ and we denote by $\epsilon_{i}$ the element in ±$\pm\{1\}$ such that $A(e_{i})=\epsilon_{i}e_{i}$. We now write $A(e_{1})=\epsilon_{1}e_{1}+\sum_{i=2}^{n}x_{i}e_{i}$ with the coefficients $x_{i}\in\mathbb{Z}$ and note that $\epsilon_{1}=\pm 1$. For every $j>1$, we have that $\alpha_{j}A\alpha_{j}^{-1}\in\widetilde{\Sigma}$. Thus $\alpha_{j}A\alpha_{j}^{-1}(e_{1})=\pm e_{1}$. It is easy to see that:
(13) $\alpha_{j}^{-1}:\begin{cases}e_{1}\mapsto 3e_{1}-2e_{j}\\ e_{j}\mapsto-e_{1}+e_{j}\\ e_{k}\mapsto e_{k}\text{ for }k\neq\{1,j\}\end{cases}$
Then one we can explicitly compute the value of $\alpha_{j}A\alpha_{j}^{-1}(e_{1})$ and see that for $i\neq\{1,j\}$, $x_{i}=0$. Since this holds for all $j$, we have that $x_{i}=0$ for all $i\geq 2$. Also, we have from the equation above that $\epsilon_{1}-\epsilon_{j}=0$ for all $j$ and hence $A=\pm I$. For $i\neq j$ and $k\neq l$, it is clear that $\sigma_{i}\sigma_{j}^{-1}$, $\alpha_{i}\alpha_{j}^{-1}$ or $\sigma_{k}\alpha_{l}^{-1}$ are not elements of $\Sigma$ and that concludes the proof of our claim. ∎
Definition 2.8.
For $k\geq 2$, we choose elements $g_{1},...,g_{k}\in\operatorname{SL}(6,\mathbb{Z})\backslash\Sigma$ as in Lemma 2.7. We define $H$ as the subgroup of $\Gamma$ generated by ±$\pm I\in\Sigma$ and the elements $g_{i}^{-1}A_{i}g_{i}$ for $i=1,...,k$ and $A_{0}$ as in Definition 2.5. We note here that $H\simeq\mathbb{Z}/2\mathbb{Z}\times\mathbb{F}_{k+1}$.
We now state some preliminaries about reduced words in such an amalgamated free product $\Gamma=\Gamma_{1}*_{\Sigma}(\Sigma\times\Lambda)$.
Definition 2.9.
Any word of the form $a_{0}h_{1}a_{1}h_{2}a_{2}\operatorname{\cdot\cdot\cdot}h_{n}a_{n}$ with $n\geq 1$ where $a_{i}\in\Lambda$ with $a_{j}\neq e$ for $j=1,...,n-1$ and $h_{i}\in\Gamma_{1}\backslash\Sigma$ for all $i$ is called reduced expression in $\Gamma$.
Remark 2.10.
We note that every element of $\Gamma\backslash(\Sigma\times\Lambda)$ can be written in this form. However such expressions are not unique. In fact two such expressions $a_{0}h_{1}a_{1}\operatorname{\cdot\cdot\cdot}h_{n}a_{n}$ and $b_{0}k_{1}b_{1}\operatorname{\cdot\cdot\cdot}k_{m}b_{m}$ define the same element in $\Gamma$ if and only if $n=m$, $a_{i}=b_{i}$ for all $i$ and there exists $\sigma_{1},...,\sigma_{n-1}\in\Sigma$ such that:
  1. (1)
    $h_{1}=k_{1}\sigma_{1}$
  2. (2)
    $\sigma_{i-1}h_{i}=k_{i}\sigma_{i}$ for $i\in\{2,...,n-1\}$
  3. (3)
    $\sigma_{n-1}h_{n}=k_{n}$
This automatically forces $b_{0}k_{1}b_{1}\operatorname{\cdot\cdot\cdot}k_{i}\sigma_{i}=a_{0}h_{1}a_{1}\operatorname{\cdot\cdot\cdot}h_{i}$ in $\Gamma$ for all $i\in\{1,...,n-1\}$. For an element $g\in\Gamma$, we define the length $|g|$ as the length of a reduced expression of $g$. By convention, we say that an element in $\Sigma$ has length 0. Comparing reduced expressions of the elements in $g^{-1}Hg$ for $g\in\Gamma$, we get the following result:
Lemma 2.11.
For any element $g\in\Gamma$, we have that $g\Sigma g^{-1}\cap H\subseteq\{\pm I\}$.
Proof.
Recall that $H$ is generated by {±$\{\pm I\}$ and the elements $g_{i}A_{i}g_{i}^{-1}$. Since we can choose $g_{i}$’s such that $g_{i}g_{j}^{-1}\notin\Sigma$, a reduced expression of an element in $H$ has length at least 3. Thus we have that $H$ does not intersect $\Sigma$ except at {±$\{\pm I\}$. Now suppose we have $g=\sigma\lambda\in\Sigma\times\Lambda$ and $h\in H$ such that $h\notin\{\pm I\}$. Then $ghg^{-1}\in\Sigma$ if and only if $\lambda h\lambda^{-1}\in\Sigma$. Once again $\lambda h\lambda^{-1}$ has a reduced expression of length at least 5 and hence it does not belong to $\Sigma$. Now let $g\notin\Sigma\times\Lambda$ and let $g=a_{0}h_{1}a_{1}...h_{n}a_{n}$ be a reduced expression for $g$ with $h_{i}\in\Gamma\backslash\Sigma$ and $a_{i}\in\Lambda$. Without loss of generality (up to inverses of letters), a reduced expression of $h$ is of the form ±$\pm g_{i_{1}}A_{i_{1}}g_{i_{1}}^{-1}...g_{i_{k}}A_{i_{k}}g_{i_{k}}^{-1}$ for indices $i_{1},i_{2},...,i_{k}$. If $a_{n}\neq e$ then the reduced expression of $ghg^{-1}$ clearly has length equal to $2|g|+|h|$ and hence it cannot be in $\Sigma$. Now suppose $a_{n}\neq e$, if $h_{n}g_{i_{1}}\in\Sigma$. So we are left with the case when $a_{n}=e$ and $g=a_{0}h_{1}a_{1}...h_{n}$.
[…]
We now prove the following ad-hoc lemma about automorphisms of such amalgamated free product groups that we need for our proof.
Lemma 2.12.
Suppose $\Gamma_{i},\Gamma^{\prime}_{i}$ for $i\in\{1,2\}$ are discrete groups with property $(T)$. Further suppose $\Sigma\neq\Gamma_{i}$ and $\Sigma^{\prime}\neq\Gamma^{\prime}_{i}$ for $i=1,2$ and $\delta^{\prime}:\Gamma=\Gamma_{1}*_{\Sigma}\Gamma_{2}\rightarrow\Gamma^{\prime}_{1}*_{\Sigma^{\prime}}\Gamma^{\prime}_{2}=\Gamma^{\prime}$ is an isomorphism between the amalgamated free product groups. Then $\delta^{\prime}$ is the composition of an inner automorphism and $\delta$ where $\delta$ satisfies exactly one of the following:
  1. (1)
    $\delta(\Gamma_{1})=\Gamma^{\prime}_{1}$, $\delta(\Gamma_{2})=\Gamma^{\prime}_{2}$ and $\delta(\Sigma)=\Sigma^{\prime}$.
  2. (2)
    $\delta(\Gamma_{1})=\Gamma^{\prime}_{2}$, $\delta(\Gamma_{2})=\Gamma^{\prime}_{1}$ and $\delta(\Sigma)=\Sigma^{\prime}$.
Proof.
As in the proof of Part 4 in [DV11, Theorem 3.1] for a subgroup $G<\Gamma$, let $|G|\coloneqq\sup\{|g|,\;g\in G\}$. We have that $|G|<\infty$ if and only if there exists $g\in G$ such that $gGg^{-1}\subset\Gamma_{1}$ or $gGg^{-1}\subset\Gamma_{2}$. Also as in [DV11, Theorem 3.1], if $G$ has property (T), then $|G|<\infty$. Applying this to the group $\delta(\Gamma_{2})$, suppose that we have a $g\in\Gamma^{\prime}$ such that $\operatorname{Ad}(g)\circ\delta(\Gamma_{2})\subset\Gamma^{\prime}_{2}$. Then let $\delta_{1}=\operatorname{Ad}(g)\circ\delta$ and note that $\delta_{1}(\Gamma_{2})\subset\Gamma_{2}$. Since we have that $\Sigma^{\prime}\neq\Gamma^{\prime}_{1}$ and $\Sigma^{\prime}\neq\Gamma_{2}$, conjugates of $\Gamma^{\prime}_{1}$ or $\Gamma^{\prime}_{2}$ do not generate the whole group $\Gamma^{\prime}$. Since $\Gamma_{1}$ is a property (T) group, there exists $h\in\Gamma^{\prime}$ such that $\operatorname{Ad}(h)\circ\delta_{1}(\Gamma_{1})\subset\Gamma^{\prime}_{1}$. Now notice that $\delta_{1}(\Gamma_{1})$ and $\delta_{2}(\Gamma_{2})$ generates $\Gamma^{\prime}$ and hence every element of $\Gamma_{1}^{\prime}$ has to be generated by elements of $h^{-1}\Gamma^{\prime}_{1}h$ and $\Gamma^{\prime}_{2}$. This implies that such an element $h$ must be of the form $kl$ where $k\in\Gamma^{\prime}_{1}$ and $l\in\Gamma^{\prime}_{2}$. Now letting $\delta_{2}=\operatorname{Ad}(l)\circ\delta_{1}$, we have that $\delta_{2}(\Gamma_{1})\subset\Gamma^{\prime}_{1}$ and $\delta_{2}(\Gamma_{2})\subset\Gamma^{\prime}_{2}$. Since $\delta_{2}$ is surjective, we have that $\delta_{2}(\Gamma_{1})=\Gamma^{\prime}_{1}$ and $\delta_{2}(\Gamma_{2})=\Gamma^{\prime}_{2}$. It follows that $\delta(\Sigma)=\Sigma^{\prime}$. On the other hand if we started with a $g\in\Gamma^{\prime}$ such that $\operatorname{Ad}(g)\circ\delta(\Gamma_{2})\subset\Gamma_{1}$, then similarly we would have an isomorphism $\delta_{2}$ such that $\delta_{2}(\Gamma_{1})=\Gamma^{\prime}_{2}$, $\delta_{2}(\Gamma_{2})=\Gamma^{\prime}_{2}$ and $\delta(\Sigma)=\Sigma^{\prime}$. ∎
Lemma 2.13.
Let $\Gamma$ as in Definition 2.1 and $\delta\in\operatorname{Aut}(\Gamma)$. Assume that $h\in\Gamma$ and $H_{0}<H$ is a finite index subgroup such that $h\delta(H_{0})h^{-1}\subset H$. Then $\delta$ is inner.
Proof.
By Lemma 2.12, up to an inner automorphism, $\delta$ is of the form $\delta_{1}*_{\Sigma}\delta_{2}$ where $\delta_{1}\in\operatorname{Aut}(\Gamma_{1})$, $\delta_{2}\in\operatorname{Aut}(\Gamma_{2})$ and $\delta(\Sigma)=\Sigma$. Since $A_{0}\in H$ and $H_{0}<H$ is a finite index subgroup, we can choose an integer $s>0$ such that $A_{0}^{s}\in H_{0}$. Let $h=a_{0}h_{1}a_{1}\operatorname{\cdot\cdot\cdot}h_{n}a_{n}$ be a reduced expression for $h$. Then we have that
(14) $h\delta(A_{0}^{s})h^{-1}=a_{0}h_{1}a_{1}\operatorname{\cdot\cdot\cdot}a_{n-1}h_{n}(a_{n}\delta_{2}(A_{0}^{s})a_{n}^{-1})h_{n}^{-1}a_{n-1}^{-1}\operatorname{\cdot\cdot\cdot}h_{1}^{-1}a_{0}^{-1}$
is a reduced expression for $h\delta(A_{0}^{s})h^{-1}$ and belongs to $H$. Note that reduced expressions in $H$ are precisely the words of the form $b_{1}b_{2}\operatorname{\cdot\cdot\cdot}b_{k}$ where each $b_{i}$ is a block containing either a letter $A_{0}^{m}$ for some integer $m$ or a block containing three letters $g_{i}A_{i}^{n}g_{i}^{-1}$ for some integer $n$. Thus we have that $a_{n}\delta_{2}(A_{0}^{s})a_{n}^{-1}$ must be of the form $A_{i}^{r_{i}}$ for $0\leq i\leq k$. As in the proof of Lemma 2.2, since $A_{i}^{r_{i}}$ for $i\neq 0$ and $\delta_{2}(A_{0}^{s})$ have distinct eigenvalues, we have that $a_{n}\delta_{2}(A_{0}^{s})a_{n}^{-1}$ has to be of the form $A_{0}^{s}$. Hence equation 14 can be written as:
(15) $h\delta(A_{0}^{s})h^{-1}=a_{0}h_{1}a_{1}\operatorname{\cdot\cdot\cdot}a_{n-1}h_{n}A_{0}^{\pm r}h_{n}^{-1}a_{n-1}^{-1}\operatorname{\cdot\cdot\cdot}h_{1}^{-1}a_{0}^{-1}$
Now comparing equation 15 with a reduced expression in $H$ and using Remark 2.10, we see that there exists $\sigma\in\Sigma$ such that $a_{0}h_{1}a_{1}...a_{n-1}h_{n}\sigma\in H$. We can conclude that $h\in H(\Sigma\times\Lambda)$. Multiplying $h$ on the left with an element of $H$, we may assume that $h\in\Sigma\times\Lambda$. Now we write $h=a\sigma^{-1}$ with $a\in\Lambda$ and $\sigma\in\Sigma$. For all $i\in\{1,...,k\}$, since $g_{i}^{-1}A_{i}g_{i}\in H$, we can find $s_{i}>0$ such that $g_{i}^{-1}A_{i}^{s_{i}}g_{i}\in H_{0}$. Thus we have that $a\sigma^{-1}\delta_{1}(g_{i})^{-1}\delta_{2}(A_{i}^{s_{i}})\delta_{1}(g_{i})\sigma a^{-1}\in H$. Note that $a(\sigma^{-1}\delta_{1}(g_{i}^{-1})\delta_{2}(A_{i}^{s_{i}})(\delta_{1}(g_{i})\sigma)a^{-1}$ is a reduced expression. Once again we compare this to reduced expressions in $H$ and conclude that either $a=e$ or $a$ is a nonzero power of $A_{0}$. Thus we have that $a\in H$. Multiplying $h$ on the left with $a^{-1}$ we may assume that $h=\sigma^{-1}\in\Sigma$. Since for all $i\in\{1,...,k\}$, the elements $g_{i}^{-1}A_{i}g_{i}A_{0}$ belong to $H$, a non zero power of these belong to $H_{0}$. Thus we have:
(16) $(\sigma^{-1}\delta_{1}(g_{i})^{-1})\delta_{2}(A_{i})\delta_{1}(g_{i})\delta_{2}(A_{0})\sigma\in H$
Once again comparing equation 16 to a standard reduced expression in $H$, and looking at the eigenvalues, we get that $\delta_{2}(A_{i})=A_{i}^{\pm 1}$ for all $i\in\{0,...,k\}$. Now by Lemma 2.6, we have that $\delta_{2}=\operatorname{id}$.
Now we note that $g_{i}A_{i}^{s_{i}}g_{i}\in H_{0}$ for all $i\in\{1,...,k\}$ and we finally get that
(17) $\sigma^{-1}\delta_{1}(g_{i})^{-1}A_{i}^{s_{i}}\delta_{1}(g_{i})\sigma\in H$
Once again from equation 17 and Remark 2.10, we find elements $\sigma_{i}\in\Sigma$ such that $\delta_{1}(g_{i})\sigma=\sigma_{i}g_{i}$ for all $i$. By Proposition 2.4, $\delta_{1}=\operatorname{Ad}(X)$ with $X\in\widetilde{\Sigma}$. Thus we have that $Xg_{i}X^{-1}\sigma=\sigma_{i}g_{i}$ and hence $X^{-1}\sigma=g_{i}^{-1}X^{-1}\sigma_{i}g_{i}$. Since $X^{-1}\sigma_{i}\in\widetilde{\Sigma}$, we have that $X^{-1}\sigma\in\bigcap_{i=1}^{k}g_{i}^{-1}\widetilde{\Sigma}g_{i}$. Thus by Lemma 2.7, we have that $X^{-1}\sigma\in\{\pm I\}$. Hence we have $X=\sigma$ or $X=-\sigma$ and in both cases, $\operatorname{Ad}(X)=\operatorname{Ad}(\sigma)$, thus concluding our proof. ∎
Lemma 2.14.
If $h\in\Gamma\backslash H$, then $hHh^{-1}\cap H$ has infinite index in $H$.
Proof.
If $hHh^{-1}\cap H$ has finite index, then taking $\delta=\operatorname{id}$ in Lemma 2.13, we have a contradiction. Hence the result follows. ∎
Lemma 2.15.
Every element $g\in\Gamma\backslash\{\pm I\}$ has an infinite conjugacy class.
Proof.
Recall that $\Gamma=\Gamma_{1}\ast_{\Sigma}(\Sigma\times\Lambda)$. It is well known that $\operatorname{PSL}(6,\mathbb{Z})$ is an icc group and hence for $g\in\Sigma$ such that $g\neq\{\pm I\}$, $g$ has an infinite conjugacy class in $\Gamma$. If $g\in(\Sigma\times\Lambda)\backslash\Sigma$, then we can pick infinitely many distinct elements $c_{k}\in\Gamma_{1}\backslash\Sigma$ and note that $c_{k}gc_{k}^{-1}$ are all distinct elements in $\Gamma$. Hence it has an infinite conjugacy class. Otherwise, consider a reduced expression $g=a_{0}b_{1}a_{1}...b_{n}a_{n}$ with $a_{i}\in\Lambda$ and $b_{i}\in\Gamma_{1}\backslash\Sigma$ as in Remark 2.10. If $a_{0}$ and $a_{n}$ are both nontrivial then we can pick infinitely many distinct $c_{k}\in\Gamma_{1}\backslash\Sigma$, and if they are both trivial we pick infinitely many distinct $c_{k}\in\Lambda$. If $a_{0}=e$ and $a_{n}\neq e$, we can pick infinitely many distinct elements $c_{k}\in\Lambda$ such that $c_{k}\neq a_{n}^{-1}$ for any $k$, and similarly for $a_{0}\neq e$ and $a_{n}=e$. In all these cases $c_{k}^{-1}gc_{k}$ are all distinct and hence $g$ has an infinite conjugacy class. ∎
2.2. Terminology on group actions
We begin this section with the notion of dissipative and conservative actions.
Definition 2.16.
Let $G\curvearrowright(X,\mu)$ be an essentially free nonsingular action of a countable group $G$ on a standard probability space $(X,\mu)$. Up to measure zero, there is a unique partition of $X$ in $G$-invariant Borel sets $D$ and $C$ such that the following holds.
  • The action $G\curvearrowright D$ admits a fundamental domain: there is a Borel set $D_{0}\subset D$ such that the sets $g\cdot D_{0}$, $g\in G$, form a partition of $D$.
  • The action $G\curvearrowright C$ is recurrent: for every non-null Borel set $C_{0}\subset C$, there are infinitely many $g\in G$ such that $\mu(C_{0}\cap g\cdot C_{0})>0$.
The action is called dissipative if $\mu(C)=0$ and is called conservative or recurrent if $\mu(D)=0$.
For the results in the later sections, we shall use the notion of induced actions which we define now:
Definition 2.17.
Let $G\curvearrowright(X,\mu)$ be a nonsingular action of a countable group and let $H<G$ be a subgroup. The action is called induced from H if there exists a $G$-equivariant Borel map $X\rightarrow G/H$.
Note that every action is trivially induced from itself. If the action $G\curvearrowright(X,\mu)$ admits a fundamental domain $E$ then it is induced from $\{e\}\curvearrowright E$. We shall use the following result proved by Schmidt and Walters in [SW82] heavily in the later sections, so we state it here for convenience. As in [AIM21, Theorem 7.3], we note that the recurrence assumption is weaker than the proper ergodicity assumption in [SW82] but the same proof holds for recurrent actions.
Theorem 2.18 (c.f. Theorem 7.3 in [AIM21]).
Let $G$ be a countable discrete group. Let $\sigma:G\curvearrowright(X,\mu)$ be a nonsingular action and $\rho:G\curvearrowright(Y,\nu)$ be a pmp action. Suppose that $\sigma$ is recurrent and $\rho$ is mixing, then every $(\sigma\times\rho)$-invariant function in $L^{\infty}(X\times Y,\mu\times\nu)$ is contained in $L^{\infty}(X)$.
We end this section by defining the notion of s-malleable actions as introduced in [Pop07].
Definition 2.19.
Let $G$ be a locally compact group and let $G\curvearrowright(X,\mu)$ be a finite or infinite measure preserving action. The action is called s-malleable if there exists a one-parameter group $(\alpha_{t})_{t\in\mathbb{R}}$ of measure preserving transformations of $X\times X$ and a transformation $\beta=\beta^{-1}$ of $X\times X$ such that the following hold:
  1. (1)
    $\alpha_{t}$ and $\beta$ commute with the diagonal action $G\curvearrowright X\times X$
  2. (2)
    $\alpha_{1}(x,y)\in\{y\}\times X$ for a.e. $(x,y)\in X\times X$
  3. (3)
    $\beta(x,y)\in\{x\}\times X$ for a.e. $(x,y)\in X\times X$
  4. (4)
    $\alpha_{t}\circ\beta=\beta\circ\alpha_{-t}$ for all $t\in\mathbb{R}$.
3. Construction
Definition 3.1.
Let $\Gamma_{1}$, $\Sigma$, $\Lambda$, $\Gamma$ and $\pi$ be as in Definition 2.1. Let $\Gamma\curvearrowright(X,\mu)$ be the generalized Bernoulli action $(X_{0},\mu_{0})^{\Gamma/H}$ where $H$ is the subgroup of $\Gamma$ as in Definition 2.8, and $X_{0}=\{0,1\}$ with $\mu_{0}(0)\neq\mu_{0}(1)$. Next we let $\Gamma_{1}\curvearrowright Y$ be the translation action of $\operatorname{SL}(6,\mathbb{Z})$ on $\mathbb{R}^{6}$. Now let $\alpha$ be the infinite measure preserving action of $\Gamma$ on $Z=X\times Y=(X_{0},\mu_{0})^{\Gamma/H}\times\mathbb{R}^{6}$ given by $g\cdot(x,y)=(gx,\pi(g)y)$. We let $\mathcal{R}$ be the type $\textrm{II}_{\infty}$ orbit equivalence relation of the action $\alpha$.
Definition 3.2.
Let $Y_{1}=\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ and consider the translation action of $\operatorname{SL}(6,\mathbb{Z})$ on $Y^{1}$. Now we let $\alpha_{1}$ be the action of $\Gamma$ on $Z_{1}=X\times Y_{1}=(X_{0},\mu_{0})^{\Gamma/H}\times\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ given by $g\cdot(x,y)=(gx,\pi(g)y)$. We let $\mathcal{R}_{1}$ be the orbit equivalence relation of the action $\alpha_{1}$.
It is possible to identify the action $\alpha_{1}$ with a different action to make computations easier. This has been done in [Ver22, Section 5], and we summarize the results in the following lemma.
Lemma 3.3.
[…]
Proof.
We first identify $\mathbb{R}^{*}_{+}$ with $\mathbb{R}$ and consider the action of $\operatorname{SL}(6,\mathbb{Z})\times\mathbb{R}$ on $\mathbb{R}^{6}$ given by:
(18) $(g,t)\cdot y=(e^{-t/6}g)y$
where the $e^{-t/6}g$ acts on a vector $y$ by matrix multiplication. Since the determinant of $g$ is 1, we have that $\det(e^{-t/6}g)=e^{-t}$. Hence the $\mathbb{R}$ action scales the measure and the $\operatorname{SL}(6,\mathbb{Z})$ action is measure-preserving. By [Ver22, Proposition 4.1.14], the quotient $\mathbb{R}^{6}/\mathbb{R}$ can be equipped with a measure with respect to which the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}/\mathbb{R}$ is nonsingular. Now we can identify the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ with $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}/\mathbb{R}$. Once again by [Ver22, Proposition 4.1.14], the Maharam extension of this action can be identified with the translation action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}$. Since the generalized Bernoulli action in Definition 3.2 is p.m.p., we have that the Maharam extension of $\mathcal{R}_{1}$ is $\mathcal{R}$. In particular since $\mathcal{R}$ is ergodic by Lemma 4.2, we have that $\mathcal{R}_{1}$ is of type $\textrm{III}_{1}$. From the identification above, the ergodicity of $\mathcal{R}$ implies that also $\mathcal{R}_{1}$ is ergodic and that concludes the proof. ∎
4. $\mathcal{U}_{\operatorname{fin}}$-cocycle superrigidity
To compute the outer automorphism group that we want, the first step is to show that the constructed action is cocycle superrigid in the following sense.
Definition 4.1.
A non-singular essentially free ergodic action $G\curvearrowright(X,\mu)$ is called cocycle superrigid with target group $H$ if any 1-cocycle $\omega:G\times X\rightarrow H$ is cohomologous to a cocycle $(g,x)\mapsto\delta(g)$ for a.e. $x\in X$ where $\delta:G\rightarrow H$ is a group homomorphism. We say that such an action is $\mathcal{U}_{\operatorname{fin}}$-cocycle superrigid if it is cocycle superrigid with any target group $H$ in Popa’s class $\mathcal{U}_{\operatorname{fin}}$, i.e. any closed subgroup of the unitary group $\mathcal{U}(M)$ for a $\textrm{II}_{1}$ factor $M$ with separable predual.
Lemma 4.2.
Let $(Z,\eta)=(X\times Y,\mu\times\nu)$ and $\Gamma\curvearrowright(Z,\eta)$ be as in Definition 3.1. Then the four-fold diagonal action $\Gamma\curvearrowright Z\times Z\times Z\times Z$ is ergodic.
Proof.
We first notice that for an element $g\in\Gamma$, the subgroup $g\Sigma g^{-1}$ does not intersect $H$ except at {±$\{\pm I\}$. Therefore the action $\Sigma\curvearrowright X$ is mixing. Since the diagonal product of mixing actions is still mixing, we have that $\Sigma\curvearrowright X\times X\times X\times X$ is still mixing. By [PV11, Lemma 5.6], we have that $\Sigma\curvearrowright Y\times Y\times Y\times Y$ is ergodic, and hence it is a conservative nonsingular action. By Theorem 2.18, we have that $\Sigma\curvearrowright Z\times Z\times Z\times Z$ is ergodic and hence the four-fold diagonal action of $\Gamma$ is also ergodic. ∎
Lemma 4.3.
The action $\Gamma\curvearrowright(Z,\eta)=(X\times Y,\mu\times\nu)$ as in Definition 3.1 is $\mathcal{U}_{\operatorname{fin}}$-cocycle superrigid.
Proof.
Let $G$ be a group in Popa’s class $\mathcal{U}_{\operatorname{fin}}$ and let $\omega:\Gamma\times(X\times Y)\rightarrow G$ be a 1-cocycle. We first consider the restriction of $\omega$ to $\operatorname{SL}(6,\mathbb{Z})\times(X\times Y)$ and show that the restriction is cohomologous to a group homomorphism, essentially using [PV11, Theorem 5.3]. We know that generalized Bernoulli actions are s-malleable and as in the proof of [PV11, Theorem 1.3], $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}$ is s-malleable. Hence the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright X\times Y$ is s-malleable. We claim that the diagonal action $\Gamma\curvearrowright X\times Y\times X\times Y$ has property (T).
We show this in the same way as in Step 1 of the proof of [KS13, Theorem 21]. Let $\mathcal{G}=SL(6,\mathbb{R})$ and consider the action $\mathcal{G}\curvearrowright\mathbb{R}^{6}$. Denoting by $e_{1},e_{2},...,e_{6}$ the basis vectors of $\mathbb{R}^{6}$, notice that the orbit of $(e_{1},e_{2})$ in the diagonal action $\mathcal{G}\curvearrowright Y\times Y$ has a measure zero complement. Let us denote by $H_{0}$ the stabilizer of the point $(e_{1},e_{2})$ under the $\mathcal{G}$ action. Then we can identify the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright(X\times Y\times X\times Y)$ with the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\operatorname{SL}(6,\mathbb{R})/H_{0}\times X\times X$. By [PV11, Proposition 3.3], $\operatorname{SL}(6,\mathbb{Z})\curvearrowright X\times Y\times X\times Y$ has property (T) if and only if $H_{0}\curvearrowright\operatorname{SL}(6,\mathbb{R})/\operatorname{SL}(6,\mathbb{Z})\times X\times X$ has property (T). Since this is a p.m.p free ergodic action, this has property (T) as $H_{0}$ has property (T) by [PV11, Proposition 3.1]. By Lemma 4.2, all the hypotheses of [PV11, Theorem 5.3] are satisfied and using the theorem, we have that $\omega|_{\operatorname{SL}(6,\mathbb{Z})}$ is cohomologous to a group homomorphism.
As in the proof of Lemma 4.2, the diagonal action $\Sigma\curvearrowright X\times Y\times X\times Y$ is ergodic. Since $\Sigma$ commutes with $\Lambda$, we have that $\omega$ is a group homomorphism on $\Lambda$ by [PV11, Lemma 5.5]. Since $\Gamma$ is generated by $\operatorname{SL}(6,\mathbb{Z})$ and $\Lambda$, we have that $\omega$ is given by a group homomorphism on $\Gamma$, and this concludes the proof. ∎
In the final proposition of this section, we shall compute the cohomology group $H^{1}(\mathcal{R}_{1},S^{1})$, i.e., the group of 1-cocycles of $\mathcal{R}_{1}$ with values in the circle up to coboundaries.
Proposition 4.4.
Let $\mathcal{R}_{1}$ be as in Definition 3.2. Then $H^{1}(\mathcal{R}_{1},S^{1})=\mathbb{R}$
Proof.
It is a known fact that $\Gamma_{1}=\operatorname{SL}(6,\mathbb{Z})$ and $\Lambda=\operatorname{SL}(3,\mathbb{Z})$ are both equal to their commutator subgroups and are hence perfect. Hence we have that $\Gamma$ is also perfect and there are no nontrivial group homomorphism from $\Gamma$ to $S^{1}$. Now suppose $\Omega:\Gamma\times Z_{1}\rightarrow S^{1}$ is a 1-cocycle. Since by Lemma 4.2, the diagonal action $\Gamma\curvearrowright Z_{1}\times Z_{1}$ is ergodic and by Lemma 4.3, the Maharam extension $\mathcal{R}$ is cocycle superrigid with target $S^{1}$, we note that all the hypotheses of [Ver22, Theorem 5.3.1] are satisfied. Now by [Ver22, Theorem 5.3.1], there exists a group homomorphism $\gamma:\mathbb{R}\rightarrow S^{1}$ such that $\Omega$ is cohomologous to the 1-cocycle:
(19) $\Gamma\times Z_{1}\ni(g,z)\mapsto\gamma(-\omega(g,z))\in S^{1}$
where $\omega$ is the logarithm of the Radon Nikodym 1-cocycle. Since the group of homomorphisms from $\mathbb{R}$ to $S^{1}$ is $\mathbb{R}$ and since each such homomorphism defines a 1-cocycle as above, we conclude that $H^{1}(\mathcal{R}_{1},S^{1})=\mathbb{R}$. ∎
5. Computation of the outer automorphism group
We begin this section with the following lemma to show that the action restricted to non-trivial normal subgroups of $\Gamma$ cannot be dissipative.
Lemma 5.1.
If $N\triangleleft\Gamma$ is a normal subgroup and $N\curvearrowright\mathbb{R}^{6}\times(X_{0},\mu_{0})^{\Gamma/H}$ is dissipative, then $N=\{I\}$ or $N=\{\pm I\}$.
Proof.
Consider the normal subgroup $N_{0}=N\cap\ker(\pi)$. Thus $N_{0}$ acts only on the second component $N_{0}\curvearrowright(X,\mu)=(X_{0},\mu_{0})^{\Gamma/H}$. Any essentially free pmp action of an infinite group is recurrent. We can choose a probability measure on $\mathbb{R}^{6}$ which is equivalent to the Lebesgue measure and note that the action $N_{0}\curvearrowright\mathbb{R}^{6}\times(X_{0},\mu_{0})^{\Gamma/H}$ is recurrent if $N_{0}$ is infinite. Since the action is dissipative, this gives a contradiction and we get that $N_{0}$ is finite. By Lemma 2.15, we have that $N_{0}\subseteq\{\pm I\}$. However since $\pi(-I)=-I$, we have that $N_{0}=\{I\}$. Hence $\pi|_{N}:N\rightarrow\operatorname{SL}(6,\mathbb{Z})$ is injective. For an element $h\in\operatorname{SL}(3,\mathbb{Z})$ and $k\in N$, we have that $\pi(hkh^{-1}k^{-1})=0$ and hence by injectivity we have $hk=kh$. Thus $N$ commutes with $\operatorname{SL}(3,\mathbb{Z})$. Therefore we have that $N\subset\Sigma\subset\operatorname{SL}(6,\mathbb{Z})$. By normality we have that for all $i$, $g_{i}^{-1}Ng_{i}=N$. Hence by Lemma 2.7 we have:
(20) $N=\bigcap_{i}g_{i}^{-1}Ng_{i}\subseteq\bigcap_{i}g_{i}^{-1}\widetilde{\Sigma}g_{i}\subseteq\{\pm I\}$
and that concludes the proof. ∎
Cocycle superrigidity often leads to orbit equivalence superrigidity results. The first step in this direction in the p.m.p setting was taken by Zimmer in [Zim20, Proposition 2.4] and then generalized by Popa in [Pop07, Theorem 5.6]. The most general result in the setting of nonsingular actions of locally compact groups was proven by Popa and Vaes in [PV11, Lemma 5.10]. We use here the following simplified version which is essentially proven in [DV23, Lemma 2.4].
Lemma 5.2.
Let $\Gamma\curvearrowright X$ and $G\curvearrowright Y$ be nonsingular essentially free ergodic actions and $\Delta:X\rightarrow Z$ be an orbit equivalence. Suppose the associated Zimmer 1-cocycle is cohomologous to a group homomorphism. Then there exists:
  1. (1)
    A normal subgroup $N\triangleleft\Gamma$ such that $N\curvearrowright X$ admits a fundamental domain.
  2. (2)
    A subgroup $G_{0}<G$ and a non-null set $Z_{0}\subset Z$ such that $Z_{0}$ is $G_{0}$-invariant and $G\curvearrowright Z$ is induced from $G_{0}\curvearrowright Z_{0}$.
  3. (3)
    A nonsingular isomorphism $\Delta_{0}:X/N\rightarrow Z_{0}$ and a group isomorphism $\delta:\Gamma/N\rightarrow G_{0}$ such that $\Delta_{0}(gx)=\delta(g)\cdot\Delta_{0}(x)$ a.e. and $\Delta(x)\in G\cdot\Delta_{0}(Nx)$ for a.e. $x\in X$.
Proposition 5.3.
Let $\Gamma\curvearrowright(Z,\eta)=(X\times Y,\mu\times\nu)$ be as in Definition 3.1 and let $\mathcal{R}$ be the orbit equivalence relation. Then for any orbit equivalence $\Delta\in\operatorname{Aut}(\mathcal{R})$, up to multiplication by an inner automorphism of $\mathcal{R}$, we have that $\Delta(g\cdot z)=g\cdot\Delta(z)$ for a.e. $z\in Z$ and for all $g\in\Gamma$.
Proof.
Let $\Delta\in\operatorname{Out}(\mathcal{R})$ be an orbit equivalence for the action. By Lemma 4.3, the associated Zimmer 1-cocycle is cohomologous to a group homomorphism. Hence we can apply Lemma 5.2 to find a normal subgroup $N$, a subgroup $\Gamma_{0}$ and a $\Gamma_{0}$-invariant subset $Z_{0}\subset X\times Y$ such that $\Gamma\curvearrowright X\times Y$ is induced from $\Gamma_{0}\curvearrowright Z_{0}$. Now by Lemma 4.2, the action is doubly ergodic, and hence by [PV11, Lemma 6.1], $\Gamma\curvearrowright X\times Y$ is not nontrivially induced, and we can assume $\Gamma_{0}=\Gamma$ and $Z_{0}=X\times Y$ up to measure zero. Since $N\curvearrowright X\times Y$ admits a fundamental domain, it is dissipative. By Lemma 5.1, $N$ is either $\{I\}$ or {±$\{\pm I\}$. We show next that $N$ can never be {±$\{\pm I\}$.
Suppose $N=\{\pm I\}$ and then as in Lemma 5.2 there must be an isomorphism $\delta:\Gamma\rightarrow\Gamma/N$. Now it can be checked that the center of $\Gamma$ is {±$\{\pm I\}$. Indeed notice that any element in the center of $\Gamma$ must be in $\Sigma$ and must commute with $\operatorname{SL}(6,\mathbb{Z})$, however it is a known fact that the center of $\operatorname{SL}(6,\mathbb{Z})$ is {±$\{\pm I\}$ and hence the center of $\Gamma$ is {±$\{\pm 1\}$. However the center of $\Gamma/N$ is then clearly trivial, hence giving us a contradiction. Hence $N=\{I\}$ and after composing $\delta$ with an inner automorphism, we have by Lemma 2.12 that $\delta:\Gamma\rightarrow\Gamma$ is of the form $\delta_{1}*_{\Sigma}(\delta_{1}|_{\Sigma}\times\delta_{2})$ where $\delta_{1}\in\operatorname{Aut}(\Gamma_{1})$ and $\delta_{2}\in\operatorname{Aut}(\Lambda)$. By point 3 of Lemma 5.2, we also have now that there is an orbit equivalence $\Delta_{0}\in\operatorname{Aut}(\mathcal{R})$ such that $\Delta_{0}=\phi\circ\Delta$ for some element ϕ[]italic-$\phi\in[\mathcal{R}]$ with the property that $\Delta_{0}(g\cdot x)=\delta(g)\cdot\Delta_{0}(x)$ for a.e. $x\in X$ and all $g\in G$.
Let us denote by $H_{1}<H$ the subgroup of index 2 generated by $g_{i}^{-1}A_{i}g_{i}$ for all $i\in\{1,...,k\}$ and $A_{0}$. Clearly $H_{1}\subset\ker(\pi)$ and since $\pi\circ\delta=\delta_{1}\circ\pi$, we also have that $\delta(H_{1})\subset\ker(\pi)$. By Lemma 2.14, we have that $H_{1}$ acts on $\Gamma/H-\{H\}$ with infinite orbits. By [PV08, Proposition 2.3], the generalized Bernoulli action $H_{1}\curvearrowright X$ is ergodic. Consider now an $H_{1}$-invariant function in $L^{\infty}(X\times Y)$. Since $H_{1}\subset\ker(\pi)$ we have that:
(21) $L^{\infty}(X\times Y)^{H_{1}}=L^{\infty}(X_{0}^{eH}\times Y)$
Now we claim that there exists $h\in\Gamma$ and a finite index subgroup $H_{0}<H_{1}$ such that $h\delta(H_{0})h^{-1}\subset H$. Suppose this does not hold, then we have that:
(22) $L^{\infty}(X\times Y)^{\delta(H)}\subset 1\otimes L^{\infty}(Y)$
From the two previous equations, we then have that:
(23) $\Delta_{0}(L^{\infty}(X_{0}^{eH}\times Y))\subset 1\otimes L^{\infty}(Y)$
Since we have that $\Delta_{0}(g\cdot x)=\delta(g)\cdot\Delta_{0}(x)$ for all $g\in\Gamma$, it follows that:
(24) $\Delta_{0}(L^{\infty}(X\times Y))\subset 1\otimes L^{\infty}(Y)$
which is absurd and thus we have proved our claim. Now by Lemma 2.13, we have that $\delta$ is inner. Suppose $\delta(g)=kgk^{-1}$ for some $k\in\Gamma$. Then by replacing $\Delta$ by $k^{-1}\Delta$, we have that $\Delta(g\cdot z)=g\cdot\Delta(z)$ for all $g\in\Gamma$ and a.e. $z\in Z$
Proposition 5.4.
Let $\Gamma\curvearrowright(Z,\eta)=(X\times Y,\mu\times\nu)$ be as in Definition 3.1. Let $\Delta\in\operatorname{Aut}(\mathcal{R})$ be an orbit equivalence satisfying $\Delta(g\cdot z)=g\cdot\Delta(z)$ for a.e. $z\in Z$ and for all $g\in\Gamma$. Then $\Delta=(\operatorname{id}\times\rho_{\lambda})$ for some $-\0\lambda\in\mathbb{R}\backslash\{0\}$ where $\rho_{\lambda}(y)=\lambda y$.
Proof.
We show that for an orbit equivalence $\Delta$ satisfying the hypotheses of the proposition, $\Delta=\operatorname{id}$ in the following three steps.
Claim 1: $\Delta(1\otimes L^{\infty}(Y))=1\otimes L^{\infty}(Y)$. Since $\ker(\pi)$ acts on $\Gamma/H$ with infinite orbits, we have that the generalized Bernoulli action of $\ker(\pi)$ on $X$ is ergodic. Now let $1\otimes f\in 1\otimes L^{\infty}(Y)$ and consider the function $g=\Delta(1\otimes f)$. We have that $g(hx,y)=g(x,y)$ for all $h\in\ker(\pi)$. For each $y\in Y$, consider the functions $g_{y}(x)=g(x,y)$. By ergodicity we have that $g_{y}$ is constant almost everywhere. Thus $g\in 1\otimes L^{\infty}(Y)$.
Claim 2: After composing $\Delta$ with $(\operatorname{id}\times\rho_{\lambda})$ for $\0\lambda\in\mathbb{R}\backslash\{0\}$ if needed, we have that $\Delta(1\otimes f)=1\otimes f$ for all $f\in L^{\infty}(Y)$. To check this, we first note that by Claim 1, the restriction of $\Delta$ to $L^{\infty}(Y)$ commutes with the action $\operatorname{SL}(6,\mathbb{Z})\curvearrowright\mathbb{R}^{6}$. By [Fur08, Theorem D], then for a.e. $y\in Y$, $\Delta(y)=g\cdot y$ for some $g\in\operatorname{GL}(6,\mathbb{R})$ that commutes with all elements in $\operatorname{SL}(6,\mathbb{Z})$. Thus $g$ is of the form $\lambda I$ and $\Delta$ restricted to $L^{\infty}(Y)$ is of the form $\rho_{\lambda}$ where $\0\lambda\in\mathbb{R}\backslash\{0\}$ as required.
Claim 3: $\Delta=(\operatorname{id}\times\rho_{\lambda})$ for some $\lambda\in\mathbb{R}/\{0\}$. To check this, first let us write $\Delta(x,y)=(\Delta_{1}(x,y),\Delta_{2}(x,y))$. Suppose that as in Claim 2, $\widetilde{\Delta}=(\operatorname{id}\times\rho_{\lambda^{-1}})\circ\Delta$ restricted to $L^{\infty}(Y)$ is the identity. Since $\widetilde{\Delta}(L^{\infty}(X_{0}^{eH}\times Y))=L^{\infty}(X_{0}^{eH}\times Y)$, we have that for a.e. $y\in Y$ the map given by $\widetilde{\Delta}_{y}(x)=\widetilde{\Delta}_{1}(x,y)$ is an automorphism of $L^{\infty}(X_{0},\mu_{0})$. Since $\widetilde{\Delta}$ must be measure scaling and $\widetilde{\Delta}|_{L^{\infty}(Y)}=\operatorname{id}|_{L^{\infty}(Y)}$, we must have that $\widetilde{\Delta}$ is measure preserving. Hence for a.e. $y\in Y$, we have that $\widetilde{\Delta}_{y}$ is measure preserving, and hence for a.e. $y\in Y$, $\widetilde{\Delta}_{y}=\operatorname{id}$ as $\mu_{0}(0)\neq\mu_{0}(1)$ thus proving the claim. ∎
To prove the uniqueness of Cartan subalgebras we use a general result for certain actions of amalgamated free product groups as in [Vae14]. As observed in [Vae14, Remark 8.3], the same proof as [Vae14, Theorem 8.1] works for the following slightly more general result about uniqueness of Cartan subalgebras in this setting:
Theorem 5.5 (c.f. Theorem 8.1 and Remark 8.3 in [Vae14]).
Let $\Gamma=\Gamma_{1}*_{\Sigma}\Gamma_{2}$ be any amalgamated free product group and assume there exists $g_{1},...,g_{n}\in\Gamma$ such that $\cap_{i=1}^{n}g_{i}\Sigma g_{i}^{-1}$ is finite. Let $\Gamma\curvearrowright(Z_{1},\eta_{1})$ be any nonsingular free ergodic action with orbit equivalence relation $\mathcal{R}_{1}$. Let $\Gamma\curvearrowright(Z,\eta)$ denote the infinite measure preserving Maharam extension with orbit equivalence relation $\mathcal{R}$. Let $\Lambda_{i}<\Gamma_{i}$ be subgroups satisfying the following condition:
  • ($\ast$)
    For every Borel set $U\subset Z$ with $0<\eta(U)<\infty$, almost every $\mathcal{R}(\Lambda_{i}\curvearrowright Z)|_{U}$ equivalence class consists of infinitely many $\mathcal{R}(\Sigma\cap\Lambda_{i}\curvearrowright Z)|_{U}$ equivalence classes.
Then $L^{\infty}(Z_{1},\eta_{1})$ is the unique Cartan subalgebra up to unitary conjugacy in $L(\mathcal{R}_{1})=L^{\infty}(Z_{1},\eta_{1})\rtimes\Gamma$.
Proof.
Let $A=L^{\infty}(Z,\eta)$ and let $\tau$ be the canonical semifinite trace on $M=A\rtimes\Gamma$. Consider a projection $p\in A$ such that $0<\tau(p)<\infty$. As in the proof of [Vae14, Theorem 8.1], for $i\in\{1,2\}$, it is enough to find unitaries $u_{i},v_{i}\in p(A\times\Gamma_{i})p$ that satisfies:
(25) $E_{p(A\rtimes\Sigma)p}(u_{i})=E_{p(A\rtimes\Sigma)p}(v_{i})=E_{p(A\rtimes\Sigma)p}(u_{i}^{*}v_{i})=0$
where $E_{B}$ for a subalgebra $B\subset M$ denotes the unique faithful normal conditional expectation onto $B$. As observed in [Vae14, Remark 8.3], by [IKT09, Lemma 2.6] we can find such unitaries $u_{i},v_{i}$ in $p(A\rtimes\Lambda_{i})p$ such that:
(26) $E_{p(A\rtimes\Sigma\cap\Lambda_{i})p}(u_{i})=E_{p(A\rtimes\Sigma\cap\Lambda_{i})p}(v_{i})=E_{p(A\rtimes\Sigma\cap\Lambda_{i})p}(u_{i}^{*}v_{i})=0$
Now notice that $E_{p(A\rtimes\Sigma)p}(x)=E_{p(A\rtimes\Sigma\times\Lambda_{i})p}(x)$ for all $x\in p(B\rtimes\Lambda_{i})p$ and hence the unitaries $u_{i},v_{i}$ satisfy equation 25, as required. ∎
Suppose as in Theorem 5.5, we have subgroups $\Lambda_{i}<\Gamma_{i}$ such that the actions of $\Lambda_{i}$ on $Z$ are recurrent and the actions of $\Sigma_{i}=\Lambda_{i}\cap\Sigma$ on $Z$ are dissipative. Then for a fixed $i\in\{1,2\}$ the action $\Sigma_{i}\curvearrowright Z$ admits a fundamental domain $U$. For any non-null Borel subset $U_{1}\subseteq gU$ for some $g\in\Sigma_{1}$, it is then clear that any $(\Lambda_{i}\curvearrowright Z)|_{U_{1}}$ orbit contains infinitely many $(\Sigma_{i}\curvearrowright Z)|_{U_{1}}$ orbits. Now writing any non-null Borel subset $E\subset Z$ with finite measure as $E=\sqcup_{g\in\Sigma_{1}}gU\cap E$, we can see that each $(\Lambda_{i}\curvearrowright Z)|_{E}$ orbit contains infinitely many $(\Sigma_{i}\curvearrowright Z)|_{E}$ orbits, thus satisfying condition ()($\ast$)( ∗ ). This gives the following corollary to Theorem 5.5:
Corollary 5.6.
Let $\Gamma=\Gamma_{1}*_{\Sigma}\Gamma_{2}$ be any amalgamated free product group and assume there exists $g_{1},...,g_{n}\in\Gamma$ such that $\cap_{i=1}^{n}g_{i}\Sigma g_{i}^{-1}$ is finite. Let $\Gamma\curvearrowright(Z_{1},\eta_{1})$ be any nonsingular free ergodic action with Maharam extension $\Gamma\curvearrowright(Z,\eta)$. Let $\Lambda_{i}<\Gamma_{i}$ be subgroups such that the actions $\Lambda_{i}\cap\Sigma\curvearrowright Z$ are dissipative and the actions $\Lambda_{i}\curvearrowright Z$ are recurrent. Then $L^{\infty}(Z_{1},\eta_{1})$ is the unique Cartan subalgebra up to unitary conjugacy in $L^{\infty}(Z_{1},\eta_{1})\rtimes\Gamma$.
Lemma 5.7.
Let $\Gamma$ be as in Definition 2.1 and $\Gamma\curvearrowright(Z_{1},\eta_{1})$ as in Definition 3.2. Then the Maharam extension $\Gamma\curvearrowright(Z,\eta)$ as in Definition 3.1 satisfies Condition ()normal-∗($\ast$)( ∗ ) of Theorem 5.5.
Proof.
Consider the subgroup $\Lambda_{1}\subset\Gamma_{1}=\operatorname{SL}(6,\mathbb{Z})$ given by a copy of $\operatorname{SL}(2,\mathbb{Z})$ in the first two rows and columns. Then we have that $\Sigma_{1}=\Lambda_{1}\cap\Sigma$ is given by:
(27) $\Sigma_{1}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}\epsilon_{1}&\mathbb{Z}\\ 0&\epsilon_{2}\end{matrix}&0\\ \hline\cr 0&I\end{array}\right)\text{ where }\epsilon_{1}=\epsilon_{2}=1\text{or }-1$
It can be checked that the action of the top left $2\times 2$ block on $\mathbb{R}^{2}$ is dissipative and hence the action of $\Sigma_{1}\curvearrowright Z$ is dissipative. Now we note that $\Lambda_{1}\curvearrowright\mathbb{R}^{2}$ is ergodic and look at the action on $Z$ as $\Lambda_{1}\curvearrowright(X\times\mathbb{R}^{4})\times\mathbb{R}^{2}$. Choose a probability measure $\omega$ equivalent to the Lebesgue measure on $\mathbb{R}^{4}$ and note that the action $\Lambda_{1}\curvearrowright(X\times\mathbb{R}^{4},\mu\times\omega)$ is still pmp. It follows from [AIM21, Theorem A.29] that the product of a recurrent action and a pmp action remains recurrent and hence the action $\Lambda_{1}\curvearrowright Z$ is recurrent. Now let $\Lambda_{2}=\operatorname{SL}(3,\mathbb{Z})\subset\Gamma_{2}$ and note that $\Lambda_{2}\cap\Sigma=\{e\}$. Moreover $\Lambda_{2}$ is an infinite group with an essentially free pmp action on $Z$ and hence the action is recurrent. Now the result follows from Corollary 5.6. ∎
Now we have gathered all the ingredients to prove our main theorem. Let $\mathcal{R}_{1}$ be as in Definition 3.2 and let $M_{1}=L(\mathcal{R}_{1})$. Let $\delta:\mathbb{R}\rightarrow\operatorname{Out}(M)$ be the canonical homomorphism that does not depend on the choice of states.
Theorem 5.8.
[…]
Proof.
By Lemma 3.3, we have that $M_{1}$ is of type $\rm{III}_{1}$, hence $M_{1}$ is full and the modular homomorphism is injective. By Lemma 5.7 and Theorem 5.5, we have that $L^{\infty}(Z,\eta)$ is the unique Cartan subalgebra in $M_{1}$ up to unitary conjugacy. Hence we have that:
(28) $\operatorname{Out}(M_{1})=\operatorname{Out}(\mathcal{R}_{1})\ltimes H^{1}(\mathcal{R}_{1},S^{1})$
By Proposition 4.4, we have that $H^{1}(\mathcal{R}_{1},S^{1})$ is equal to $\mathbb{R}$. The only thing left to show is that $\operatorname{Out}(\mathcal{R}_{1})$ is trivial. Indeed for any orbit equivalence $\Delta_{1}\in\operatorname{Aut}(\mathcal{R}_{1})$, we claim that $\Delta_{1}$ is inner. First note that $\Delta_{1}$ induces an orbit equivalence of the Maharam extension $\mathcal{R}$, let us call it $\Delta$. By Proposition 5.3, up to an inner automorphism $\phi$ of $\mathcal{R}$, we have that $\Delta(g\cdot z)=g\cdot\Delta(z)$ for all $g\in\Gamma$ and a.e. $z\in Z$. Now by Proposition 5.4, we have that $\Delta=\phi\circ(\operatorname{id}\times\rho_{\lambda})$ for some $\0\lambda\in\mathbb{R}\backslash\{0\}$. Now, $\Delta$ restricted to $\mathcal{R}$ is equal to $\Delta_{1}$, and we have that $\Delta|_{\mathcal{R}_{1}}=\phi|_{\mathcal{R}_{1}}\circ(\operatorname{id}\times\rho_{\lambda})|_{\mathcal{R}_{1}}=\Delta_{1}$. Note that $\phi|_{\mathcal{R}_{1}}\in[\mathcal{R}_{1}]$, as for a.e. $z\in Z$ and $s\in\mathbb{R}$, we have that $((z,s),\phi(z,s))\in\mathcal{R}$ and this implies $(z,\phi|_{\mathcal{R}_{1}}(z))\in\mathcal{R}_{1}$. Let the restriction of $\rho_{\lambda}:\mathbb{R}^{6}\rightarrow\mathbb{R}^{6}$ be denoted by $\theta:\mathbb{R}^{6}/\mathbb{R}^{*}_{+}\rightarrow\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$. Note that if $\lambda>0$ then $\theta(y)=y$ for all $y\in\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$ and if $\lambda<0$, then $\theta(y)=-y$ for all $y\in\mathbb{R}^{6}/\mathbb{R}^{*}_{+}$. In both cases, $(\operatorname{id}\times\theta)\in[\mathcal{R}_{1}]$ and this proves that $\Delta\in[\mathcal{R}_{1}]$. ∎